OpenStudy (anonymous):

A cylinder of cheese is to be removed from a spherical piece of cheese with a radius of 8cm What is the maximum volume of the cylinder of cheese?

OpenStudy (anonymous):

well i hope i dont miss anything here.. but i would suggest to find a relation between the height of the cylinder (call it h) and its radius (call it r) :

OpenStudy (anonymous):



OpenStudy (anonymous):

that's good thanks :)

OpenStudy (anonymous):

well wait!!

OpenStudy (anonymous):

i wrote: now you have: r^2+8^2=(0.5h)^2 which leads to: r^2=0.25(h^2)-64 it is wrong. it should be r^2+(0.5h)^2=8^2 r^2=64-0.25(h^2)

OpenStudy (anonymous):

now you can write your function: volume-of-cylinder: pi*h*r^2 = (using the relation we got) = pi*h*(64-0.25(h^2))

OpenStudy (anonymous):

sorry for the confusion lol

OpenStudy (anonymous):

that's fine :p